# Neutron and $\bm{\gamma}$-ray Discrimination by a Pressurized Helium-4 Based Scintillation Detector

### v2

@@ 3.2 Neutron Response Systematics by G4
As mentioned in Sec. 1, neutron scatters off a 4He nucleus transferring a part of its kinetic energy ($E_{n}$). The recoiling 4He deposits the recoil energy within the fiducial volume of the detector via ionization, followed by scintillation. The recoiling 4He deposits its entire energy within the volume, which is designated as $E_{dep}$. The $E_{dep}$ spectra for monoenergetic neutrons at different energies ranging from 0.25 to 3 MeV are estimated as shown in the Fig. 9. Since there is spread in the $E_{dep}$ spectra, we have estimated the 90% value of the total area under the spectral profile. Subsequently, we have determined the upper limit of the $E_{dep}$ parameter for the integral under the spectral profile, which matches with the above-mentioned number. We have designated the corresponding upper limit on $E_{dep}$ as the marker ($\Delta E$) of the corresponding recoil signal registered by the detector. A plot of ($\Delta E$) as function of $E_{n}$ is shown in the Fig. 9. The ($\Delta E$) is found to be fairly linear with the incident energy $E_{n}$. Therefore, from the $E_{dep}$ spectrum, an estimate of the incident energy may be done.
@@
It may be noted that the recoiling 4He deposits its entire energy within the volume. Also, one neutron is found to cause multiple nuclear recoil during its passage. Therefore, each recoil is also recorded as a distinct event for systematic studies and calculation of S1 signal (to be explained in Sec. 4). The HDPE slabs are cumulatively placed and the simulations are done for 50 million events for each thickness of the absorber. Only the fast neutron events selected by energy cut of $\lesssim 70$ keV, are recorded. Fig. 10 shows the distribution of the $E_{dep}$ corresponding to the fast neutron events, which can be compared with the spectra obtained experimentally (see Fig. 6). The relative ordering of the plots for different thicknesses of the HDPE absorbers agrees with the experiment.
[…]
Figure 10: $E_{dep}$ spectrum of the (a) NR events and (b) ER events from G4 simulation. In both the cases, the neutrons from the 241Am source pass through the HDPE layers before falling on the detector.
3.3 Electron recoil events
$\gamma$-rays falling on the PHe detector cause electrons inside the gas to recoil and thus, result in electron recoil signal. This is demonstrated in our experiment (see Fig. 3) with 137Cs standard $\gamma$-ray source as well. The range of $E_{dep}$ values can be seen to merge with the lower range of $E_{dep}$ values for the neutrons as shown in the Fig. 4, where the spectra for the neutrons and the $\gamma$-rays from 252Cf are overlaid.
It has already been demonstrated by exposing the detector with 241Am source that the neutrons are attenuated by the HDPE layers (see Sec. 2.3). It is also shown that while the $\gamma$-ray spectrum at the lower channels along the $x$-axis get significantly attenuated (see Fig. 7) after passing though the HDPE layers, the higher channels corresponding the larger $E_{dep}$ values get populated as well. For measurements with different number of HDPE layers in between, we observe that after traversing two HDPE layers, the spectrum at the higher channels (i.e, region of interest or RoI $>150$) show a rising trend and a bump around 240 channel number till it reaches saturation at the highest channel. There is not much visible relative change for four HDPE layers. However, the spectral population over the RoI diminish after passing through five HDPE layers. This may be interpreted as due to the production of 2.225 MeV $\gamma$-rays from neutron moderation within the HDPE, followed by absorption in HDPE through the $n+p=d+\gamma$ process. A simulation of the $\gamma$-ray interaction with the detector is done to understand the origin, the nature of the $\gamma$-ray spectra and the underlying systematics of the process.
The $\gamma$ rays falling on the detector will produce electron recoil within the pressurized 4He-volume. The study is more complicated due to the presence of stainless steel housing of the detector which contains high $Z$ elements. Since the $\gamma$-rays are energetic to undergo pair creation, positrons are also produced, which result in the production of 511 keV $\gamma$-rays. Therefore, the electron-recoil events, though largely contain effects due to the electrons, there would be effect due to the positrons as well. Fig. 10 includes the energy loss spectra for the electron recoil events, when the neutrons emitted from the 241Am source, are incident on the detector. Henceforth, we will designate the energy loss by the electrons as $E_{dep}^{ee}$ to distinguish it from $E_{dep}$ due to the neutrons.
3.4 Gamma-ray Response Systematics by G4
In the simulation procedure, mono-energetic $\gamma$-rays of different energies ($E_{\gamma}$) are allowed to impinge upon the detector. The $E_{dep}^{ee}$ spectrum of the recoiling electrons is shown in Fig. 11. A peak begins to appear at $E_{dep}^{ee}\sim 280$ keV as the energy of the incident $\gamma$-ray is increased. The $\gamma$-ray energy threshold for appearance of the peak is found to be around 650 keV. Therefore, the 847 keV $\gamma$-ray, which arises from the ${}^{56}{\rm Fe}(n,n^{\prime}\gamma)$ reaction, would also give rise to the $\sim 280$ keV peak due to the presence of SS housing, when no HDPE slab is placed. A detailed event-by-event tracking reveals that this number corresponds to the most probable $E_{dep}^{ee}$ of the recoiling electrons given the geometry (diameter) of the detector. The $\Delta E$ values for the $E_{dep}^{ee}$ spectra as function of $E_{\gamma}$ is plotted in the Fig. 11, which indicates saturating trend at higher energies. The contour is fitted with a quadratic function, modulated by a smooth step function of the form:
$F(x)=0.5\cdot(1-\tanh((x-p0)/p1))\cdot(p5\cdot x^{2}+p3\cdot x+p2)+0.5\cdot p4\cdot(1+\tanh((x-p0)/p1))$
where, $p0$ to $p4$ are the parameters for the fit. The step-function makes the quadratic term dominant in the low $E_{\gamma}$ region and the constant term (saturation value) dominant in the high $E_{\gamma}$ region. Specifically, $p0$ marks the threshold when the saturation begins dominating, while $p4$ denotes the saturation value.
[…]
Figure 11: (a) $E_{dep}^{ee}$ spectrum of mono-energetic $\gamma$-rays, and (b) corresponding 90% contour values are plotted as function of the $\gamma$-ray energy.
The saturating behaviour in the Fig. 11 arises due to the fact that the recoiling electrons do not lose their entire energy within the detector. This can be seen more explicitly in Fig. 12, where the $E_{dep}^{ee}$ is plotted against their kinetic energy (KE) for mono-energetic incident $\gamma$-rays of 1 MeV and 2.225 MeV. The kinetic energy of the recoiling electrons increases with $E_{\gamma}$ as expected. The sharp right edge in the plots is the Compton edge. It is evident from Fig. 12 that a significant fraction of the electrons with higher recoil energies lose only $\sim$280 keV of their energy within the detector. Therefore, we can conclude that the peak observed at higher channel numbers in the Fig. 6 corresponds to $\sim 280$ keV of $E_{dep}^{ee}$.
[…]
Figure 12: $E_{dep}^{ee}$ vs. recoil kinetic energy of ER events from mono-energetic $\gamma$ sources for (a) 1.0 MeV and (b) 2.22 MeV
To further confirm this geometric effect, the simulation was repeated by changing the diameter of the fiducial volume of the detector to 30 mm and 90 mm. Corresponding $E_{dep}^{ee}$ spectra are plotted in the Fig. 13. A clear shift in the peak can be seen, which is at $\sim$130 keV for 30 mm diameter (solid line) and $\sim$450 keV for 90 mm (dotted line).
Another systematic study was done to understand the effect of the SS cover of the detector. In this case, the SS cover was removed (keeping the original diameter of 60 mm unchanged for the active volume) and the corresponding $E_{dep}^{ee}$ spectrum was obtained. It was observed that the height of the peak at $\sim$280 keV had reduced in the spectrum indicating that the peak was, indeed, enhanced by the presence of the SS cover. The contribution from pair-production events to the RoI and in the overall spectra was evaluated from the $E_{dep}^{ee}$ spectrum obtained after vetoing those events where pair-production has taken place. No significant deviation was observed from the spectrum of Fig. 11, indicating that pair-production does not make significant contribution to the spectra. Additionally, it was observed from simulation that in almost all the cases, a single recoil electron is produced from a single $\gamma$-ray. The number of events with two recoil electrons from the incident $\gamma$-ray is almost two orders of magnitude lower than the single recoil events.
Figure 13: $E_{dep}^{ee}$ spectra with 30 (solid lines), 60 (dot-dashed lines) and 90 mm (dotted lines) diameter of the SS housing.
Based on the experimental and the simulation results on the $\gamma$-ray response of the PHe detector, we can conclude that: 1) the $E_{dep}^{ee}\sim 280$ keV peak arising due to the most probable value of energy loss by the recoiling electrons within the confined geometry of the PHe detector. This is also seen as rising bump at the higher channel numbers ($\sim 240$) in the experiment (see Fig. 7), apart from the spectral feature at the lower channel numbers ( $\leq\,150$) which has been demonstrated as arising from accompanying low energy $\gamma$-rays of the so-called room background. Based on the spectral profile of the $E_{dep}^{ee}$ spectra, we fitted the peak with logistic function after subtracting an exponential background for the $E_{dep}^{ee}$ spectra shown in the Fig. 11, from 662 keV to 2.225 MeV $\gamma$-rays. The average $E_{dep}^{ee}$ value for the peak was found to be: $282\pm 10$ keV. Based on the above and the fact that the $x$-axis of the plots of the Fig. 7 scale as $E_{dep}^{ee}$, a first order calibration with calibration constant of 1.175 keV/channel on the $x$-axis is applied. 2) Observed attenuation of the low energy $\gamma$-rays by the HDPE layers, which is mostly due to scattering of the $\gamma$-rays, qualitatively agrees with the simulation. However, no quantitative comparison could be attempted in our study due to lack of information about the discrete $\gamma$-rays from the sources. 3) Production of the 2.225 MeV $\gamma$-rays and correlation with the attenuation of the fast neutrons by the HDPE layers have been found to demonstrate reasonable agreement between our experiment and the G4 simulation.
3.5 Neutron spectral calibration based on the neutron energy spectral data of 252Cf
@@
It is demonstrated through our G4 simulation that the $\Delta E$ values of the energy deposition by the neutrons inside the detector have a linear dependence of the neutron energy $E_{n}$ (see Sec. 3.2 and Fig. 9). Based on this observation, an alternative energy calibration of the detector was attempted using the prompt neutrons from a 252Cf source. It is expected that the energy spectrum of the neutrons originating from the spontaneous fission is Maxwellian. However, after detailed analysis by various groups on the properties of neutron sources, and several meetings under the banner of International Atomic Energy Agency (IAEA), held during 1980 to 1987, the expert group on 252Cf spontaneous fission-based fast neutron sources, had proposed a corrected Maxwellian spectrum, based on experimental data and theoretical estimates[17] over various energy segments from 0.2 MeV to 20 MeV. Accordingly, the corrected Maxwellian spectrum $F(E_{n})$ as function of the kinetic energy ($E_{n}$) of the neutrons for the prompt fission neutrons is given by
$F(E_{n})=R(E_{n})\,\frac{2}{\sqrt{\pi}}\,\frac{\sqrt{E_{n}}}{T^{3/2}}\exp[{-\frac{E_{n}}{T}}],$ (1)
@@
4 Scintillation signal estimates for the PHe detector
In our simulation work, we have so far followed the transfer of energy from the radiation quanta (neutrons or $\gamma$-rays) to the recoiling medium particles (4He or electron). There are primarily three mechanisms by which, a recoiling 4He (NR events) loses its energy along its track, viz. ionization, charge-exchange collisions and excitation collisions. The cross-sections of these interactions depend on the charge-state of the recoiling 4He itself. Electron-ion pairs are produced by ionization. While in charge-exchange collisions, ions are generated through electron-capture interactions by the recoiling He1+ and He2+ ions. Conversely, free electrons are produced from He0 and He1+ ions through the loss of electrons. When these electron-ion pairs recombine, Helium excimer states are formed. Additionally, they are also formed when excited He-atoms resulting from excitation collisions, combine with ground state He-atoms. Therefore, one of the main outcomes of these interactions is the formation of the Helium excimer molecules. These excimer molecules can be in spin-singlet or spin-triplet states (their proportion dependent on the production mechanism). The spin-singlet states decay promptly via radiative transitions and give rise to prompt scintillation, which is termed as the S1 signal, whereas the spin-triplet states are long-lived but they too decay radiatively and contribute to the delayed S3 signal[21].
[…]
Figure 15: (a) Yield vs. recoil energy of 4He. The e--ion pair yield is shown in blue and excitation yield is shown in red; (b) S1 and S3 signal intensities vs. recoil energy for NR (red) and ER (blue) events in pressurized 4He.
The above mechanism also works for the ER events, except for the absence of charge-exchange collisions. From the discussion, it can be concluded that the determining parameter for the intensity of these signals is the number of electron-ion pairs and excitation produced for a given recoil energy. The calculation of the yield (number per unit recoil energy) and the signal intensities have been worked out in detail [21]. The density of helium-4 gas at 180 bar pressure and 300 K temperature is taken as $0.0247\,{\rm g}\cdot{\rm cm}^{-3}$ and the number density is 3.7$\times$1021 cm-3. Figure 15 shows the yield vs. recoil energy of 4He, where the blue curve is the electron-ion ionization yield and the red curve is the excitation yield. The ionization and excitation yields due to ER are considered constant at 22.7 keV-1 and 10.2 keV-1 respectively [21]. Figure 15 depicts the S1 and S3 signal intensities as a function of the recoil energy. The blue and the red curves represent NR and ER events respectively. The plot of the S1 signal distribution for NR and ER events for incident neutrons is shown in Fig. 16 and Fig. 16 respectively for different number of the HDPE layers placed in between. The incident energy of the neutrons is sampled from the energy distribution obtained from the 11B ($\alpha,n$)14N reaction (see Fig. 7). The $\gamma$-rays produced from the 241Am source and also due to interaction of the source neutrons with the intervening media have been taken into account for the S1-S3 signal estimates.
[…]
Figure 16: S1 signal distribution for (a) NR events, and (b) ER events.
We have plotted the events on the S1-S3 plane for incident neutrons to estimate the bands corresponding to NR and ER events. This is shown in the Fig. 17. Here the narrow band towards the left side is due to the ER events (from the $\gamma$-rays produced in the nuclear reactions), while the band towards the right side is due to the NR events. In the NR band, most of the events cluster within a curve-like region (indicated by red). This is a consequence of saturation of the S1 signal at larger recoil energies of 4He, as illustrated in the Fig. 15.
Based on the above simulations, we can conclude that by combining the S1 and S3 scintillation signals arising from the fast and the slow decay of the ${}^{4}{{\rm He}^{*}}_{2}$ excimer states, a discrimination between NR and the ER events can be done. Our present simulation-based study does not have a scope to estimate the realistic background which can constrain the above threshold. Unusually long decay time ($\sim 13\,{\rm s}$) of the triplet state leading to the S3 signal may also severely limit the suitability of the proposed method.
Figure 17: S3 vs. S1 for NR and ER events due to the incident neutrons.
@@ 5 Conclusion
Pressurized 4He detector, capable of detecting and discriminating between the neutrons and the $\gamma$-rays in a mixed radiation field, has been evaluated in detail both through source-based experiment and G4-based simulation.
The experimental studies based on 252Cf fast neutron source, 137Cs $\gamma$-reference source and 241Am source have successfully demonstrated that a fast and a slow gate-based integration of the detector output pulses by the QDC is effective in achieving discrimination. Attenuation of the neutron band by the HDPE layers placed between the source and the detector also confirms the discrimination method followed. The charge contents of the pulses over the long gate effectively scale as the energy deposited $E_{dep}$ by the recoiling 4He in case of the neutrons, or the recoiling electrons in case of the $\gamma$-rays. Based on the estimated fast neutron energy spectrum arising from the spontaneous fission of 252Cf[17], and the PHe detector efficiency data[19] for the fast neutrons, the measured $E_{dep}$ spectral profile has been found to match reasonably well with theoretically estimated profile, folded by the detector efficiency. The lower energy part of the spectral profile, however, do not match due to the difference in the threshold for our measurement and the PHe detector efficiency data, which was based on the time-of-flight measurements[19]. The measured spectral distribution was calibrated over a range in terms of the neutron energy, based on the above match of the profiles. The same calibration was applied for the measured neutron spectra pertaining to the 241Am source based measurements.
@@
Interaction of the neutrons and the $\gamma$-rays with the HDPE layers was investigated in detail through Geant4 based simulation. Production of the neutrons from the 241Am source was demonstrated to be largely due to the ($\alpha,n$) reactions on the constituent elements of the glass substrate. Corresponding neutron spectrum was used for the event generation. Neutron attenuation by the HDPE layers was successfully reproduced to qualitatively match with the experimental results. Interaction of the $\gamma$-rays emitted from the 241Am source and also produced through neutron absorption by the HDPE layers with the pressurized helium active medium, were studied to find the origin of the peak-like structures in the measured spectra. It is found that a significant number of relatively higher energy $\gamma$-rays ($E_{\gamma}\gtrsim 650\,{\rm keV}$) produce recoiling electrons, which dump around 280 keV of energy, resulting in the peak-like structure. Role of the SS housing of the detector and its geometry were also investigated as part of the systematic studies in support of the above.
The scintillation signals from the pressurized 4He has been estimated following the formalism of Guo and McKinsey[21]. Estimation of the S1 and S3 signals due to scintillation were worked out. We have demonstrated that a combination of the (S1,S3) signal can be used as a potential method to discriminate between the neutrons and the $\gamma$-rays. It may be useful for a future dark matter direct search experiment, provided a very low threshold on the NR-ER discrimination can be achieved. Alternatively, a time projection chamber (TPC) based on pressurized 4He as the active medium, should be investigated where the possibility of discrimination using the scintillation (S1) and the ionization (S2) signals may be explored. As mentioned before, relatively smaller number of available electrons make the PHe detector much less sensitive to the $\gamma$-ray / electron recoil background.
[…]
Figure 18: Projected sensitivity limit at 90% CL with no background; XENONnT upper limit at 90% CL (2023) result for reference [24]
@@ 6 Acknowledgements
We are grateful to Prof. Satyaki Bhattacharya of SINP, India, for the invaluable discussions on the statistical analysis for the sensitivity calculation. We would also like to acknowledge technical and operational support from Chandranath Marick of SINP, India. One of us (SS) would like to acknowledge financial support from the Department of Atomic Energy Raja Ramanna Fellowship (DAE-RRF) scheme to carry out this work.

### v3

@@ 2.5 Neutron Energy Calibration based on 252Cf Spectral Data
As mentioned in Sec. 1, neutron scatters off a 4He nucleus transferring a part of its kinetic energy ($E_{n}$). The recoiling 4He deposits the recoil energy within the fiducial volume of the detector via ionization, followed by scintillation. The recoiling 4He deposits its entire energy within the volume, which is designated as $E_{dep}$. The $E_{dep}$ spectra for monoenergetic neutrons at different energies ranging from 0.5 to 3 MeV are estimated as shown in the Fig. 13. Since there is spread in the $E_{dep}$ spectra, we have estimated the 90% value of the total area under the spectral profile. Subsequently, we have determined the upper limit of the $E_{dep}$ parameter for the integral under the spectral profile, which matches with the above-mentioned number. We have designated the corresponding upper limit on $E_{dep}$ as the marker ($\Delta E$) of the corresponding recoil signal registered by the detector. A plot of ($\Delta E$) as function of $E_{n}$ is shown in the Fig. 13. The ($\Delta E$) is found to be fairly linear with the incident energy $E_{n}$. Therefore, an estimate of the incident energy may also be done from the $E_{dep}$ spectrum.
@@
Based on this observation, an alternative energy calibration of the detector was attempted using the prompt neutrons from a 252Cf source. It is expected that the energy spectrum of the neutrons originating from the spontaneous fission is Maxwellian. However, after detailed analysis by various groups on the properties of neutron sources, and several meetings under the banner of International Atomic Energy Agency (IAEA), held during 1980 to 1987, the expert group on 252Cf spontaneous fission-based fast neutron sources, had proposed a corrected Maxwellian spectrum, based on experimental data and theoretical estimates[16] over various energy segments from 0.2 MeV to 20 MeV. Accordingly, the corrected Maxwellian spectrum $F(E_{n})$ as function of the kinetic energy ($E_{n}$) of the neutrons for the prompt fission neutrons is given by
$F(E_{n})=R(E_{n})\,\frac{2}{\sqrt{\pi}}\,\frac{\sqrt{E_{n}}}{T^{3/2}}\exp[{-\frac{E_{n}}{T}}],$ (2.1)
@@ 3 Conclusion
PHe detector, capable of detecting and discriminating between the neutrons and the $\gamma$-rays in a mixed radiation field, has been evaluated in detail through source-based experiment and supplemented by G4-based simulation.
The experimental studies based on 252Cf fast neutron source, 137Cs $\gamma$-reference source and 241Am source have successfully demonstrated that a fast and a slow gate-based integration of the detector output pulses by the QDC is effective in achieving discrimination. Attenuation of the neutron band by the HDPE layers placed between the source and the detector also confirms the discrimination method followed. The charge contents of the pulses over the long gate effectively scale as the energy deposited $E_{dep}$ by the recoiling 4He in case of the neutrons, or the recoiling electrons in case of the $\gamma$-rays.
@@
Interaction of the neutrons and the $\gamma$-rays with the HDPE layers was investigated in detail through G4-based simulation. Production of the neutrons from the 241Am source was demonstrated to be largely due to the ($\alpha,n$) reactions on the constituent elements of the glass substrate. Corresponding neutron spectrum was used for the event generation. Interaction of the $\gamma$-rays emitted from the 241Am source and also produced through neutron absorption by the HDPE layers with the pressurized Helium active medium, were studied to find the origin of the peak-like structures in the high-end tail of the measured spectra. It is found that a significant number of relatively higher energy $\gamma$-rays ($E_{\gamma}\gtrsim 650\,{\rm keV}$) produce recoiling electrons, which dump around 280 keV of energy, resulting in the peak-like structure. This observation was used to derive a first-order calibration in terms of $E_{dep}^{ee}$, which was then applied to the measured neutron spectra. Role of the detector geometry was investigated as part of the systematic studies in support of the above.
@@ 4 Acknowledgements
We would like to acknowledge technical and operational support from Chandranath Marick of SINP, India. One of us (SS) would like to acknowledge financial support from the Department of Atomic Energy Raja Ramanna Fellowship (DAE-RRF) scheme to carry out this work.